OpenStudy (anonymous):

write the slope-intercept equation for the line with slope -2/3 and y-intercept(0,3)and graph the line

OpenStudy (anonymous):

use y-y1 = m(x-x1) y-3 = (-2/3)(x-0) y = (-2/3)x

OpenStudy (anonymous):

mr bubbles! madame, how do you do? http://www.youtube.com/watch?v=PmUnqrZI4qE

OpenStudy (anonymous):

mz. bubblez that is satelite73

OpenStudy (anonymous):

im good